# Baltimore Atlas: FreqWeaver Adapter for Semi-supervised Ultra-high Spatial Resolution Land Cover Classification

### v1

@@ 3. Method
We propose a semi-supervised, parameter-efficient fine-tuning framework built upon the Segment Anything Model 2 (SAM2), which effectively leverages the prior knowledge of large-scale foundation models. To this end, we introduce the FreqWeaver Adapter—a lightweight yet effective module that adapts SAM2 to remote sensing data while maintaining high parameter efficiency. In addition, we design a semi-supervised learning strategy that fully utilizes unlabeled data to improve domain adaptation and generalization under limited supervision. The overall framework is illustrated in Figure 1.
3.1. FreqWeaver: Frequency-Modulated Adapter for Efficient RS Fine-Tuning
Remote sensing (RS) imagery is typically characterized by significant scale variation, densely distributed objects, and complex textures. These properties pose challenges not only to general-purpose vision models such as SAM, but also to existing parameter-efficient tuning methods, which often struggle to adapt to such complex data distributions. To address this, we propose FreqWeaver, a lightweight frequency–temporal modulation adapter. FreqWeaver explicitly decomposes input features into low- and high-frequency components, which are processed separately using convolutional filters with distinct receptive fields to model global semantics and fine-grained details, respectively. In addition, a temporal branch is introduced to capture residual semantic information that may not be effectively represented in the frequency domain. This design significantly improves the adaptability and structural representation capability of SAM on RS imagery, while maintaining parameter efficiency.An overview of the proposed adapter is illustrated in Figure 2.
3.1.1. Spectral Decomposition
To facilitate frequency-aware feature refinement, we perform spectral decomposition on the encoder outputs using the 2D Discrete Fourier Transform (DFT). This allows explicit separation of low- and high-frequency information for targeted processing.
Before applying the DFT, we first reshape the input feature tensor from $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ to $\mathbf{X}^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$ to align with the channel-wise Fourier transform convention.
@@ 3.1.1. Spectral Decomposition
To extract different frequency bands, we define two complementary radial binary masks:
@@
where $r$ denotes the radial distance from the spectral center, and $\rho$ is a predefined cutoff threshold.
We isolate the low- and high-frequency components by applying the masks element-wise:
@@
Finally, we recover the spatial-domain representations via the inverse Fourier transform:
@@
resulting in real-valued tensors $\widehat{\mathbf{X}}_{\mathrm{lf}},\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ representing the low- and high-frequency spatial features, respectively. These components are subsequently passed to dedicated branches for frequency-specific refinement.
Figure 2. Overview of the proposed FreqWeaver Adapter.
3.1.2. Frequency–Temporal Feature Modulation
To enhance frequency- and spatial-domain representations in a disentangled manner, FreqWeaver introduces three specialized branches, each responsible for modeling a specific component of the feature spectrum: low-frequency (global structure), high-frequency (local detail), and spatial-domain semantics. These branches operate independently and produce complementary signals for subsequent fusion.
@@ 3.1.2. Frequency–Temporal Feature Modulation
(5) $\mathbf{F}_{\mathrm{lf}}=\mathcal{F}_{\mathrm{LF}}(\widehat{\mathbf{X}}_{\mathrm{lf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the input low-frequency feature map, and $\mathcal{F}_{\mathrm{LF}}(\cdot)$ denotes a depthwise convolution with a kernel size of $11\times 1$ 11, initialized with zeros. The output $\mathbf{F}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ encodes coarse structural information and is reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
@@
The high-frequency response is defined as:
(6) $\mathbf{F}_{\mathrm{hf}}=\mathcal{F}_{\mathrm{HF}}(\widehat{\mathbf{X}}_{\mathrm{hf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the high-frequency input, and $\mathcal{F}_{\mathrm{HF}}(\cdot)$ is a depthwise convolution with a $5\times 5$ kernel. The output $\mathbf{F}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ highlights spatially localized structures and is similarly reshaped to $\mathbb{R}^{B\times H\times W\times C}$.
@@
This branch directly processes the spatial-domain feature without spectral filtering, enabling the modeling of mid-frequency signals and ambiguous semantic cues that may not be cleanly separated in the frequency domain. It plays a complementary role in capturing transitional or occluded content.
The spatial-domain response is formulated as:
(7) $\mathbf{F}_{\mathrm{s}}=\mathcal{F}_{\mathrm{S}}(\mathbf{X}),$
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original feature map before frequency decomposition. The operator $\mathcal{F}_{\mathrm{S}}(\cdot)$ is a two-layer bottleneck MLP. The output $\mathbf{F}_{\mathrm{s}}\in\mathbb{R}^{B\times H\times W\times C}$ captures spatial-domain semantics that complement the frequency-based branches.
3.1.3. Content-Aware Feature Fusion
To integrate the outputs of the frequency–temporal branches, we adopt a content-aware fusion strategy guided by global contextual information. This mechanism allows the model to dynamically balance the relative contributions of low-frequency, high-frequency, and spatial-domain features on a per-sample basis.
We first compute a global context descriptor via global average pooling over the spatial dimensions:
@@ 3.1.3. Content-Aware Feature Fusion
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original spatial-domain feature map.
@@
(9) $[r_{\mathrm{s}},r_{\mathrm{lf}},r_{\mathrm{hf}}]=\operatorname{softmax}(\mathcal{F}_{\mathrm{Router}}(\mathbf{g}))\in\mathbb{R}^{B\times 3},$
where $\mathcal{F}_{\mathrm{Router}}(\cdot)$ denotes a two-layer MLP with softmax normalization to ensure $\sum_{k}r_{k}=1$.
Finally, the three modulation outputs are adaptively fused with the original features via weighted addition:
@@
This fusion strategy enables adaptive emphasis of complementary information sources, thereby improving the model’s ability to generalize to diverse spatial patterns in remote sensing imagery.
3.2. Uncertainty-Aware Teacher–Student Adaptation
In the context of remote sensing segmentation, acquiring large-scale annotated datasets is both expensive and labor-intensive. To alleviate this constraint, we introduce an Uncertainty-Aware Mean Teacher (UAMT) framework that leverages unlabeled data via a dual-network consistency paradigm. The method enhances pseudo-supervision reliability by integrating entropy-based uncertainty estimation into a perturbation-consistent learning scheme. This design allows for stable adaptation of large vision foundation models (e.g., SAM2) with minimal supervision.
@@ 3.2.2. Entropy-Based Uncertainty Estimation
While teacher predictions provide valuable supervisory cues, pseudo-labels generated for remote sensing images are particularly error-prone due to complex backgrounds, fine-grained object boundaries, and domain-specific variations. To mitigate the impact of such noisy supervision, we employ a pixel-wise confidence estimation mechanism based on Shannon entropy to downweight uncertain predictions.
Let $\mathbf{P}^{\text{T}}_{j}\in\mathbb{R}^{C}$ denote the softmax probability vector predicted by the teacher at pixel $j$, where $C$ is the number of classes. The uncertainty score is quantified by entropy:
@@
which attains its maximum at $\log C$ when the prediction is maximally ambiguous (i.e., uniform distribution). To normalize this measure and convert it into a usable confidence weight, we define:
@@
where higher $w_{j}$ values indicate more confident predictions and hence stronger training signals. This design enables the model to suppress the impact of high-entropy (uncertain) regions and focus on reliable pseudo-labels.
@@ 3.2.3. Loss Function: Supervised–Unsupervised Coupling
The overall objective consists of two complementary terms: a supervised loss on labeled data and a confidence-weighted consistency loss on unlabeled data.
@@
Given labeled samples $(x_{l},y_{l})\in\mathcal{D}_{l}$, the student model is trained with standard pixel-wise cross-entropy:
@@
Consistency Loss with Confidence Weighting.
For unlabeled images $x_{u}\in\mathcal{D}_{u}$, the teacher generates soft predictions $\mathbf{P}^{\text{T}}$, while the student predicts $\mathbf{P}^{\text{S}}$ from a strongly perturbed view. A pseudo-label $\hat{y}^{\text{T}}_{j}=\arg\max_{c}\mathbf{P}^{\text{T}}_{j,c}$ is assigned per pixel, and the confidence-weighted consistency loss is formulated as:
(15) $\mathcal{L}_{\mathrm{cons}}=\frac{1}{|\mathcal{D}_{u}|\cdot N}\sum_{x_{u}\in\mathcal{D}_{u}}\sum_{j=1}^{N}w_{j}\cdot\mathrm{CE}(\mathbf{P}^{\text{S}}_{j},\hat{y}^{\text{T}}_{j}),$
where $N$ is the total number of pixels per image, and $\mathbf{P}^{\text{S}}_{j}$ is the predicted softmax from the student at pixel $j$.
@@
The combined loss is given by:
(16) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{sup}}+\lambda\cdot\mathcal{L}_{\mathrm{cons}},$
where $\lambda=0.1$ balances the contributions of labeled and unlabeled samples.
During training, only the parameters of the FreqWeaver adapters are updated, while the SAM2 backbone remains frozen. This strategy ensures efficient and stable transfer to remote sensing domains, leveraging both labeled and unlabeled data in a principled manner.
Figure 3. Visualization of the State-of-the-Art PEFT Method.
@@ 4.1. Dataset
In this study, all experiments were conducted on ultra-high-resolution remote sensing imagery of Baltimore, captured at a spatial resolution of 0.3 meters. A total of 60 images, each with dimensions of 1024×1024 pixels, were selected based on the nine urban semantic categories defined by UrbanWatch, ensuring a relatively balanced distribution of class labels, as shown in Fig 4. The dataset comprises a total of 62,914,560 pixels.For initial annotations, we utilized the UrbanWatch model, which was originally trained on 1-meter resolution imagery. Although not optimized for ultra-high-resolution inputs, this model offered a coarse yet informative baseline. To improve annotation accuracy, we subsequently performed meticulous manual corrections. Given the complexity of remote sensing scenes and the difficulty of pixel-level labeling, this process was both labor-intensive and time-consuming.For training and evaluation, we used 55 images for training and the remaining 5 for testing. The test set alone contained 5,242,880 pixels, offering a sufficiently large sample size for robust and reliable performance evaluation.
@@ 4.2. Evaluation Metrics
To comprehensively evaluate the segmentation performance, we adopt two widely used metrics: Intersection over Union (IoU) and Dice coefficient. Both metrics quantify the region-level overlap between predicted segmentation results and ground truth annotations, providing a general assessment of segmentation accuracy.
@@ 4.3. Training and Testing Details
All experiments were conducted using PyTorch on a single NVIDIA RTX L40S GPU with 48 GB of memory. In our work, we adopt SAM2 as the backbone network. Input images were uniformly resized to $1024\times 1024$ during both training and testing. The model was trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 2.
For the segmentation task, we employed the standard Cross-Entropy loss, placing emphasis on accurate pixel-wise classification as the primary optimization objective.
Table 1. Comparison with state-of-the-art parameter-efficient fine-tuning (PEFT) methods.
@@
Figure 5. Visualization of the State-of-the-Art LULC Method.

### v2

@@ 3. Method
3.1. Overview
We propose Baltimore Atlas, a framework for ultra-high spatial resolution (UHSR) land cover classification. To support this framework, we construct the Baltimore Atlas Dataset, a benchmark created from 0.3 m aerial imagery of Baltimore City acquired through the National Agriculture Imagery Program (NAIP). For the model, we adopt SAM2 as the foundation and introduce the FreqWeaver Adapter, which enables parameter-efficient fine-tuning and adapts SAM2 to the multi-scale patterns of UHSR data. In addition, we design the Uncertainty-Aware Teacher–Student Framework, a semi-supervised learning method that employs uncertainty estimation to generate high-confidence pseudo-labels, which are then used as supervisory signals for training on unlabeled data. Overall, the integration of these components enables accurate land cover classification on UHSR data with limited annotated samples, as illustrated in Fig. 1.
3.2. FreqWeaver Adapter
A key challenge of UHSR imagery lies in its significant scale variation, dense object distribution, and complex textures, which limit the adaptability of SAM2 and existing parameter-efficient tuning (PEFT) methods. To address this, we introduce FreqWeaver, a lightweight frequency–spatial modulation adapter that improves multi-scale modeling while keeping parameter cost low. The method integrates frequency-domain decomposition with convolutional processing: it separates low- and high-frequency components through dedicated convolutional branches and incorporates spatial information for more complete feature representation. Our approach is organized into three stages—Frequency Decomposition, Multi-Branch Feature Modeling, and Adaptive Feature Fusion—which will be described in the following sections. An overview of the proposed adapter is illustrated in Figure 2.
3.2.1. Frequency Decomposition
To enable frequency-aware feature refinement, we apply frequency decomposition to the encoder features using the 2D Discrete Fourier Transform (DFT). This operation explicitly separates low- and high-frequency components for specialized processing.
Before applying the DFT, the input feature tensor is reshaped from $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ to $\mathbf{X}^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$ to align with the channel-wise Fourier transform convention.
@@ 3.2.1. Frequency Decomposition
To extract different frequency bands, two complementary radial binary masks are defined:
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function, $r$ is the radial distance from the spectral center, and $\rho$ is a predefined cutoff threshold.
The low- and high-frequency components are isolated by applying the masks element-wise:
@@
Finally, the spatial-domain representations are recovered via the inverse Fourier transform:
@@
yielding real-valued tensors $\widehat{\mathbf{X}}_{\mathrm{lf}},\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ that represent the low- and high-frequency spatial features. These components are then forwarded to separate branches for frequency-specific refinement.
Figure 2. Overview of the proposed FreqWeaver Adapter. The FreqWeaver Adapter is inserted into specific Transformer layers as the only trainable component, while all other parameters remain frozen. It performs frequency decoupling by applying a Fourier transform to separate high- and low-frequency components. High-frequency information is processed with small convolution kernels to capture local textures and fine details, while low-frequency information is modeled with large kernels to enhance global boundary features. To prevent information loss, the original spatial features are preserved, and a router is employed to fuse multi-scale information effectively. This design provides more comprehensive feature representations and enables the model to better adapt to the multi-scale characteristics of UHSR data.
3.2.2. Multi-Branch Feature Modeling
To enhance frequency- and spatial-domain representations in a disentangled manner, FreqWeaver introduces three specialized branches, each responsible for modeling a specific component of the feature representations: low-frequency (global structure), high-frequency (local detail), and spatial-domain semantics. These branches operate independently and produce complementary signals for subsequent fusion.
@@ 3.2.2. Multi-Branch Feature Modeling
(5) $\mathbf{F}_{\mathrm{lf}}=f_{\mathrm{lf}}(\widehat{\mathbf{X}}_{\mathrm{lf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the input low-frequency feature map, and $f_{\mathrm{lf}}(\cdot)$ denotes a depthwise convolution with a kernel size of $11\times 11$. The output $\mathbf{F}_{\mathrm{lf}}\in\mathbb{R}^{B\times C\times H\times W}$ encodes coarse structural information and is reshaped into $\mathbb{R}^{B\times H\times W\times C}$.
@@
The high-frequency response is computed as:
(6) $\mathbf{F}_{\mathrm{hf}}=f_{\mathrm{hf}}(\widehat{\mathbf{X}}_{\mathrm{hf}}),$
where $\widehat{\mathbf{X}}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ is the high-frequency input, and $f_{\mathrm{hf}}(\cdot)$ is a depthwise convolution with a $5\times 5$ kernel. The output $\mathbf{F}_{\mathrm{hf}}\in\mathbb{R}^{B\times C\times H\times W}$ highlights spatially localized structures and is reshaped into $\mathbb{R}^{B\times H\times W\times C}$.
@@
This branch directly processes the spatial-domain features without frequency filtering. It is designed to recover residual semantics that may be overlooked by the low- and high-frequency branches, such as transitional regions, occluded objects, or mixed textures. The spatial-domain response is defined as:
(7) $\mathbf{F}_{\mathrm{s}}=f_{\mathrm{s}}(\mathbf{X}),$
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ is the original feature map before frequency decomposition, and $f_{\mathrm{s}}(\cdot)$ is a two-layer bottleneck MLP. The output $\mathbf{F}_{\mathrm{s}}\in\mathbb{R}^{B\times H\times W\times C}$ provides complementary spatial semantics to the frequency-based branches.
3.2.3. Adaptive Feature Fusion
To integrate the outputs of the low-frequency, high-frequency, and spatial branches, we adopt a content-aware fusion strategy guided by global contextual information. This mechanism enables the model to dynamically balance the relative contributions of the three feature types on a per-sample basis.
We first compute a global context descriptor by performing global average pooling over the spatial dimensions:
@@ 3.2.3. Adaptive Feature Fusion
where $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ denotes the original spatial-domain feature map.
@@
(9) $r_{\mathrm{s}},r_{\mathrm{lf}},r_{\mathrm{hf}}]=\operatorname{softmax}(f_{\mathrm{router}}(\mathbf{g}))\in\mathbb{R}^{B\times 3},$
where $f_{\mathrm{router}}(\cdot)$ denotes a two-layer MLP with softmax normalization to ensure $\sum_{k}r_{k}=1$.
Finally, the three branch outputs are fused with the original features via weighted addition:
@@
This fusion strategy adaptively balances complementary information sources, thereby improving the model’s ability to generalize to diverse spatial patterns in remote sensing imagery.
3.3. Uncertainty-Aware Teacher–Student Framework
To address the noise in pseudo-labels for semi-supervised remote sensing segmentation, especially in UHSR imagery, we propose the Uncertainty-Aware Mean Teacher (UAMT) framework. The method employs entropy-based uncertainty estimation to select high-confidence pseudo-labels as supervisory signals for unlabeled data, reducing the impact of uncertain predictions. By combining this mechanism with a perturbation-consistent teacher–student scheme, UAMT stabilizes pseudo-supervision and enables efficient adaptation of SAM2 under limited annotation.
@@ 3.3.2. Entropy-Based Uncertainty Estimation
Pseudo-labels in remote sensing segmentation are often noisy due to complex backgrounds, fine boundaries, and domain-specific variations. To reduce their negative impact, we apply pixel-wise confidence weighting based on Shannon entropy.
Let $\mathbf{P}^{\text{T}}_{j}\in\mathbb{R}^{C}$ be the teacher’s softmax probability vector at pixel $j$, where $C$ is the number of classes. The uncertainty is measured by entropy:
@@
which reaches $\log C$ when the distribution is uniform. We normalize this score and convert it to a confidence weight:
@@
where higher $w_{j}$ values indicate more reliable predictions and thus stronger supervision. This scheme downweights uncertain regions and emphasizes high-confidence pseudo-labels.
@@ 3.3.3. Loss Function: Supervised–Unsupervised Coupling
The training objective combines a supervised loss on labeled data and a confidence-weighted consistency loss on unlabeled data.
@@
For labeled pairs $(x_{l},y_{l})\in\mathcal{D}_{l}$, the student is trained with pixel-wise cross-entropy:
@@
Consistency Loss.
For unlabeled inputs $x_{u}\in\mathcal{D}_{u}$, the teacher produces probabilities $\mathbf{P}^{\text{T}}$, from which a pseudo-label $\hat{y}^{\text{T}}_{j}=\arg\max_{c}\mathbf{P}^{\text{T}}_{j,c}$ is obtained per pixel. The student predicts $\mathbf{P}^{\text{S}}$ from a perturbed view. Using confidence weights $w_{j}$ derived from entropy, the consistency loss is:
(15) $\mathcal{L}_{\mathrm{cons}}=\frac{1}{B\cdot N}\sum_{x_{u}\in\mathcal{D}_{u}}\sum_{j=1}^{N}w_{j}\cdot\mathrm{CE}(\mathbf{P}^{\text{S}}_{j},\hat{y}^{\text{T}}_{j}),$
where $B$ is the batch size and $N$ is the number of pixels per image.
@@
The final loss is:
(16) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{sup}}+\lambda\mathcal{L}_{\mathrm{cons}},$
with $\lambda=0.1$ controlling the balance between labeled and unlabeled data.
During training, only the parameters of FreqWeaver adapters are updated, while the SAM2 backbone remains frozen, ensuring efficient transfer to remote sensing domains.
@@ 4.1. Dataset
In this study, we introduce the Baltimore Atlas Dataset, an ultra-high-resolution remote sensing segmentation dataset constructed from 0.3 m aerial imagery of Baltimore City acquired through the National Agriculture Imagery Program (NAIP). The imagery is annotated into nine urban semantic categories defined by UrbanWatch through a two-stage process. In the first stage, we use the UrbanWatch model, trained on 1 m resolution imagery, to generate the initial annotations as a coarse baseline. In the second stage, these annotations are refined through manual corrections to improve accuracy. For sampling, we extract 1024 × 1024 pixel patches (300 m × 300 m) from the entire imagery to meet SAM2’s input requirements, while applying uniform sampling to maintain class balance. This process yields 60 samples (62,914,560 pixels in total), with the class distribution shown in Fig. 3. Among them, 55 are used for training and 5 for testing, providing enough scale and diversity for effective model training and reliable performance evaluation.
@@
Figure 4. Visualization of the State-of-the-Art PEFT Method. In the first row, our method shows better capture of textures and boundaries of small objects in dense urban areas, such as residential swimming pools. In the second and third rows, parking lots are recognized at the level of individual vehicles. Although this differs from region-level annotations, it highlights the strength of our local detail modeling and suggests directions for future fine-grained classification.
@@ 4.2. Evaluation Metrics
To comprehensively evaluate the segmentation performance, we adopt two widely used metrics: Intersection over Union (IoU) and Dice coefficient. Both metrics quantify the region-level overlap between predicted segmentation results and ground truth annotations, but with different emphases: IoU focuses on overall overlap consistency, while Dice is more sensitive to small objects and class imbalance. In combination, these metrics provide a comprehensive and complementary assessment of segmentation accuracy.
@@ 4.3. Training and Testing Details
All experiments are conducted using PyTorch on a single NVIDIA RTX L40S GPU with 48 GB of memory. In our work, we adopt SAM2 as the backbone network. Input images are uniformly resized to $1024\times 1024$ during both training and testing. The model is trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 2.
For the segmentation task, we employ the standard Cross-Entropy loss, placing emphasis on accurate pixel-wise classification as the primary optimization objective.
Table 1. Quantitative Comparison with State-of-the-Art Parameter-Efficient Fine-Tuning (PEFT) Methods
@@
Figure 5. Visualization of the State-of-the-Art LULC Method. The results show that earlier methods do not work well on ultra-high spatial resolution data: (1) they have trouble outlining building boundaries, for example, nearby structures are often blurred together; (2) they miss small objects, such as residential swimming pools; and (3) they mix up similar classes, for instance confusing roads with parking lots. In contrast, our method gives clear improvements on all these points.